# CHARME: A chain-based reinforcement learning approach for the minor embedding problem

### v1

@@ 3.2. Graph representation and the structure of the policy
In this section, we present an efficient representation of logical and hardware graphs and introduce a GNN-based structure for the models representing the policy.
@@
As we mentioned in the former section, the logical graph $P$ at step $t$ is presented by a tuple $(X_{P}^{(a)},X_{P}^{(f)})$. In this work, we use the constant feature matrix $X_{P}^{(f)}=[1]^{|V_{P}|\times 1}$ for all logical graph $P$. On the other hand, the hardware graph $H$ presented by a tuple $(X_{H}^{(a)},X_{H}^{(f)})$. Specifically, the feature of a node $u$ in $H$ corresponds to the logical node embedded in $u$. If node $u$ is unembedded, the feature of $u$ is $-1$.
Based on the representation of $P$ and $H$, we propose a GNN-based structure for models representing the RL-Agent’s policy (Figure 4). Specifically, this structure is applied for the actor and the critic in the RL-Agent. The structure receives an input as a tuple of $(P,H,\phi)$ and returns the action probability (for the actor) and state value (for the critic). The flow of the structure starts with passing the representations of $P$ and $H$ through GNN layers. GNN layers are powerful for learning and analysing graph-structured data, so they are capable of extracting information from the logical and hardware graphs which is essential for the minor embedding problem. Subsequently, given the numbers of GNN output features corresponding to $P$ and $H$ as $d^{\prime}_{P}$ and $d^{\prime}_{H}$ respectively, we obtain updated feature vectors $Y_{P}\in\mathbb{R}^{|V_{P}|\times d^{\prime}_{P}}$ and $Y_{H}\in\mathbb{R}^{|V_{H}|\times d^{\prime}_{H}}$ which are aggregation of topological information and initial features.
(8) $Y_{P}=GNN(X_{P}^{(a)},X_{P}^{(f)})$
(9) $Y_{H}=GNN(X_{H}^{(a)},X_{H}^{(f)})$
@@
We then group nodes in $H$ that belong to the same chain in $\phi$ and aggregate their features. In order to accomplish that, we present $\phi$ as a matrix $M_{\phi}\in\{0,1\}^{|V_{H}|\times|V_{P}|}$ such that for $\forall v\in V_{P},u\in V_{H}$, $M_{\phi}[u,v]=1$ if and only if $u\in\phi(v)$. As a result, we obtain the aggregated features $Y_{aggr}\in\mathbb{R}^{|V_{P}|\times d^{\prime}_{H}}$ as follows:
(10) $Y_{aggr}=Y_{H}\times M_{\phi}$
@@
This step is important for several reasons. First of all, it eliminates unembedded nodes in the hardware graph. The information related to unembedded nodes is considered redundant because, given a fixed hardware topology, the position of unembedded nodes can be inferred from the embedded nodes. Hence, the information of embedded nodes alone is sufficient to make the decision about the next action. As a result, filtering out unembedded nodes significantly reduces computational costs without hindering the convergence of the policy’s parameters. On the other hand, the information of nodes embedded in the same chain is aggregated. After the aggregation, the output serves as the features of chains in the hardware graph.
@@
We observe that the final embedding feature $Y_{final}$, which combines the information from the logical graph $P$, the hardware graph $H$, and the corresponding embedding $\phi$, has $O(|V_{P}|)$ dimensions. Thus, our proposed architecture addresses the challenge of the state processing by efficiently combining the information from the graphs and the embedding.
@@ 13             $\phi$:= TOPOLOGY ADAPTING$(H,\phi)$
We now present a state transition algorithm that computes the next state based on the current state and action. This algorithm integrates with the NODE EMBEDDING and TOPOLOGY ADAPTING algorithms introduced in (Ngo et al., 2023a).
Algorithm 1, referred to as STATE TRANSITION algorithm, takes current state $s_{t}$ at the time step $t$ including current logical graph $P$, current hardware graph $H$ and current embedding $\phi^{(t)}$, as well as an action $a_{t}$ as inputs. It produces the next state $s_{t+1}$ including the logical graph $P$, the hardware graph $H$ and next embedding $\phi^{(t+1)}$.
Before explaining this algorithm in details, we recall the set of embedded node as $U^{(t)}_{P}$ which can be inferred from $\phi^{(t)}$. Algorithm 1 must guarantee that the resulting next embedding $\phi^{(t+1)}$ is a feasible embedding of $P[U^{(t)}_{P}\cup\{a_{t}\}]$ into $H$. To accomplish that, in details, first we initialize the decision variable $isolated$ as True (line 2) and assign $\phi$ as $\phi^{(t)}$ (line 3). Then, the node $a_{t}$ is embedded into the current hardware graph $H$ by finding an additional embedding $\phi^{\prime}$ through NODE EMBEDDING algorithm. If the additional embedding $\phi^{\prime}$ is not empty ($a_{t}$ is successfully embedded), Algorithm 1 constructs new embedding $\phi^{(t+1)}$ by combining $\phi$ and $\phi^{\prime}$ (line 7). Besides, it updates new features for the hardware graph $H$ following a hardware representation which we discuss in the previous section (line 8). The next state $s_{t+1}$ is a combination of logical graph $P$, hardware graph $H$ with updated features, and new embedding $\phi^{(t+1)}$. In the end, we assign the value of $isolated$ as False, finish the loop and return the new state $s_{t+1}$. Otherwise, if $\phi^{\prime}=\emptyset$, we cannot embed the node $a_{t}$ with current embedding. We call this phenomenon as isolated problem. In order to overcome this bottleneck, we expand current embedding $\phi$ by TOPOLOGY ADAPTING algorithm (line 11), and repeat the whole process until we find a satisfied additional embedding $\phi^{\prime}$.
From the lemmas provided in the work (Ngo et al., 2023a), it can be proven that assuming the current embedding $\phi^{(t)}$ is feasible, the next embedding $\phi^{(t+1)}$ of the subsequent state $s_{t+1}$ returned by Algorithm 1 is a feasible embedding of $P[U^{(t)}_{P}\cup\{a\}]$ in $H$. Additionally, the empty embedding $\phi^{(0)}$ initialized at step 0 is a feasible embedding of the empty logical graph $P[\emptyset]$ in $H$. Consequently, considering $T=|V_{P}|$, we can recursively infer that the final embedding $\phi^{(T)}$ corresponding to the terminal state $s_{T}$ obtained in the $T$-th step is a feasible embedding of $P$ in $H$. Hence, with the state transition algorithm, our CHARME effectively addresses the challenge of ensuring feasibility.
@@ 3.4. Exploration Strategy
Input: Set of generated logical graphs with $m$ elements $\mathcal{G}=\left(P_{1},\ldots,P_{m}\right)$, the hardware graph $H=(V_{H},E_{H})$, the set of rescaling constants =$\mathcal{B}=\{\beta_{1},\dots,\beta_{m}\}$, the sampling limit $D$ and the exploration limit $K$
@@
2Initialize $\mathcal{O}=\left(O_{1},\ldots,O_{m}\right)$ where each $O_{i}$ is a random permutation of the set of nodes $V_{i}$
3Initialize return score $\mu_{i}=0$ for each graph $P_{i}$ and store in =$\mathcal{M}=\left\{\mu_{1},\mu_{2},\mu_{3},\ldots,\mu_{m}\right\}$
@@
7      Calculate $\zeta_{i}$ as the efficiency score of $O_{i}$
8      $msg,O^{\prime},\_:=\operatorname{ORDER\_REFINING}(P_{i},H,0,\varnothing,K,\zeta_{i})$
9      if $msg$ is True then
@@ 10             $O_{i}\leftarrow O^{\prime}$
11            Calculate $\zeta_{i}$ as the efficiency score of $O_{i}$
12            $\mu_{i}=\frac{\zeta_{i}}{\beta_{i}}$
13
14
@@ 12            $\mu_{i}=\frac{\zeta_{i}}{\beta_{i}}$
A challenge our proposed RL model needs to address is its efficiency when training with large instances. Specifically, as the number of nodes in the training logical graphs increases, the number of possible action sequences grows exponentially. Consequently, it becomes more challenging for the model to explore good action sequences for learning, potentially leading to an inefficient policy. In CHARME, we handle this issue by incorporating an exploration term in the reward function as Equation 7. The philosophy behind the exploration term is that at beginning training steps, when the policy is still developing, the model aims to replicate precomputed actions $\bar{a}_{t}$ from an existing method. Subsequently, as the policy undergoes sufficient updates, the model gradually reduces its reliance on precomputed actions and begins to explore by itself. By doing that, the model can jump into a region with good solutions at the beginning and start to explore from that region, making the training more efficient.
Precomputed actions can be extracted from the solutions of the embedding method in the work (Ngo et al., 2023a). However, that method focuses on providing a fast minor embedding, so the sequence of precomputed actions are not carefully refined. In this section, we propose a technique to explore efficient sequences of precomputed actions in training logical graphs, making the training process more efficient. To avoid confusion, from here, we mention the sequences including precomputed actions as embedding orders. We name our proposed technique as Order Exploration.
@@ 3.4.1. Overview
Given a logical graph $P=(V_{P},E_{P})$ with $T=|V_{P}|$ and a hardware graph $H=(V_{H},E_{H})$, an embedding order $O=(\bar{a}_{1},\dots,\bar{a}_{T})$ with $\bar{a}_{t}\in V_{P},t\in[1,T]$ results in a set of embeddings $\{\phi^{(1)},\dots,\phi^{(T)}\}$. As mentioned in previous sections, given an embedding step $t$, the embedding $\phi^{(t)}$ is constructed by inputting Algorithm 1 with $\phi^{(t-1)}$ and $\bar{a}_{t-1}$ for $t\in[1,T-1]$. We note that the construction of $\phi^{(t)}$ may require the topology expansion (i.e. the function $TOPOLOGY\_ADAPTING$ is triggered). We call embeddings that require topology expansion as expansion embeddings. Otherwise, we call them as non-expansion embeddings. The expansion and non-expansion notions are used later in this section. In addition, we denote $F(O)=|\phi^{(T)}|$ as the efficiency score of $O$. Order Exploration aims to fast find emebedding orders for training logical graphs such that their efficiency scores are as small as possible.
@@
The overview of Order Exploration is described in Algorithm 2. Algorithm 2 takes the inputs as the set of $m$ given training logical graphs, denoted as $\mathcal{G}=(P_{1},\ldots,P_{m})$, the set of rescaling constants =$\mathcal{B}=\{\beta_{1},\dots,\beta_{m}\}$, the sampling limit $D$ and the exploration limit $K$. The training graph $P_{i}$ includes the set of nodes $V_{i}$ and the set of edges $E_{i}$. Given a logical graph $P_{i}\in\mathcal{G}$, we define the embedding order $O_{i}=(\bar{a}_{1},...,\bar{a}_{|V_{i}|})$ as the sequence of precomputed actions for the training graph $P_{i}$. We can see that the position of an action in $O$ is the step in which the action is embedded in the hardware graph, so from here, we mention the index of actions in $O$ as the embedding step. The goal of Algorithm 2 is to fast explore $m$ embedding orders for $m$ logical training graphs, denoted as $\mathcal{O}=(O_{1},\ldots,O_{m})$, such that their efficiency scores are as small as possible.
@@
In general, Algorithm 2 randomly picks a graph $P_{i}\in\mathcal{G}$ and explore the order for $P_{i}$. In details, first, we initialize embedding orders corresponding to training graphs in $\mathcal{G}$ (line 1). Specifically, given a graph $P_{i}=(V_{i},E_{i})$, we initialize $O_{i}$ as a random permutation of the set $V_{i}$. Then, we assign each graph $P_{i}$ with an equal potential score $\mu_{i}$ (line 2). These scores indicate the potential of being further improved for embedding orders.
@@
After initialization, we start to update embedding orders within $D$ steps. For each step, a graph $P_{i}$ is sampled from the set $\mathcal{G}$ based on the distribution for potential scores (line 4). Specifically, the selection probability is given as follow:
(11) $p\left(P_{i}\mid\mathcal{M}\right)=\frac{\mu_{i}}{\sum_{i}\mu_{i}}$
Afterward, we calculate the efficiency score $\zeta_{i}$ of $O_{i}$ (line 5). Then, we refine the order $O_{i}$ by the subroutine Order Refining which is discussed later (line 6). In general, Order Refining aims to find an order $O^{\prime}$ such that its efficiency score is lower than $\zeta_{i}$ in $K$ exploration steps. The number $K$ controls the running time of this subroutine. In addition, Order Refining returns $msg$ as $True$ if a better order $O^{\prime}$ is found (line 7). If so, we update $O_{i}$ as $O^{\prime}$ (line 8), recalculate the efficiency score $\zeta_{i}$ and update the potential score $\mu_{i}$ as follows:
(12) $\mu_{i}=\frac{\zeta_{i}}{{\beta_{i}}}$
@@
Here, we can select the rescaling constant $\beta_{i}$ as the total number of qubits used for embedding $P_{i}$ resulting from an existing embedding method. Thus, the score $\mu_{i}$, calculated by Equation 12, give us the gap between our refined solution and the existing solution. Consequently, the larger the gap is, the higher the likelihood that the corresponding graph will be selected for further refinement. It is also worth noting that if the initial potential scores are sufficiently large, every graph can be selected at least once.
When the algorithm finishes discovering, it returns all feasible orders through the set $\mathcal{O}$ (line 11).

### v2

@@ 3.2. Graph Representation and the Structure of the Policy
In this section, we present an efficient representation of logical and hardware graphs and introduce a GCN-based structure for the models representing the policy.
@@
As mentioned in the previous section, the logical graph $P$ at step $t$ is presented by a tuple $(X_{P}^{(a)},X_{P}^{(f)})$. In this work, we use the constant feature matrix $X_{P}^{(f)}=[1]^{|V_{P}|\times 1}$ for all logical graphs $P$. On the other hand, the hardware graph $H$ presented by a tuple $(X_{H}^{(a)},X_{H}^{(f)})$. Specifically, the feature of a node $u$ in $H$ corresponds to the logical node embedded in $u$. If node $u$ is unembedded, the feature of $u$ is 1-1.
Based on the representation of $P$ and $H$, we propose a GCN-based structure for the models that represent the RL-Agent policy (Figure 3). Specifically, this structure is applied to the actor and the critic in the RL-Agent. The structure receives an input as a tuple of $(P,H,\phi)$ and returns the action probability of the actor and the state value of the critic. An important component of our proposed structure is a multilayer Graph Convolutional Network (GCN) (Kipf and Welling, 2017) which is powerful for learning and analyzing graph-structured data (Henaff et al., 2015; Niepert et al., 2016). In GCNs, the inputted graph is specified by an adjacency matrix $X^{(a)}$ and a feature matrix $X^{(f)}$. Those inputted matrices are propagated through a stack of $L$ layers represented by matrices $H^{(0)},\dots,H^{(L)}$ by the layer-wise propagation rule as follows:
$H^{(l+1)}=ReLU(\hat{D}^{\frac{-1}{2}}\hat{X}^{(a)}\hat{D}^{\frac{-1}{2}}H^{(l+1)}W^{(l)})$
Here, we have the adjacency matrix with added self-connections $\hat{X}^{(a)}=X^{(a)}+I$, the diagonal matrix of degrees $\hat{D}$ where $\hat{D}_{ii}=\sum_{j}\hat{X}^{(a)}_{ij}$, the trainable weight matrix of the l-th layer $W^{(l)}$, hland the ReLU activation function. The propagation rule is applied iteratively from $H^{(1)}$ to $H^{(L)}$ with the first layer $H^{(0)}$ initialized as $X^{(f)}$. For simplification, we denote this information passing mechanism as a non-linear function named $GCN$. Specifically, $GCN$ takes the adjacency matrix and the feature matrix as inputs and returns $H^{(L)}$ as output.
The flow of the structure starts with passing the representations of $P$ and $H$ through $GCN$. Subsequently, we obtain output matrices $Y_{P}\in\mathbb{R}^{|V_{P}|\times d^{\prime}_{P}}$ and $Y_{H}\in\mathbb{R}^{|V_{H}|\times d^{\prime}_{H}}$ respectively as follows:
(10) $Y_{P}=GCN(X_{P}^{(a)},X_{P}^{(f)})$
(11) $Y_{H}=GCN(X_{H}^{(a)},X_{H}^{(f)})$
@@
We then group the nodes in $H$ that belong to the same chain in $\phi$ and aggregate their features. To do this, we present $\phi$ as a matrix $M_{\phi}\in\{0,1\}^{|V_{H}|\times|V_{P}|}$ such that for $\forall v\in V_{P},u\in V_{H}$, $M_{\phi}[u,v]=1$ if and only if $u\in\phi(v)$. As a result, we obtain the aggregated features $Y_{aggr}\in\mathbb{R}^{|V_{P}|\times d^{\prime}_{H}}$ as follows:
(12) $Y_{aggr}=Y_{H}M_{\phi}$
@@
This step is important for several reasons. First, it eliminates unembedded nodes in the hardware graph. Information related to unembedded nodes is considered redundant because, given a fixed hardware topology, the position of unembedded nodes can be inferred from the embedded nodes. Hence, the information from the embedded nodes alone is sufficient to make a decision about the next action. As a result, filtering out unembedded nodes significantly reduces computational costs without hindering the convergence of the policy parameters. On the other hand, the information of nodes embedded in the same chain is aggregated. After aggregation, the output serves as the features of the chains in the hardware graph.
@@
We observe that the final embedding feature $Y_{final}$, which combines the information from the logical graph $P$, the hardware graph $H$, and the corresponding embedding $\phi$, has $O(|V_{P}|)$ dimensions. Thus, our proposed architecture addresses the challenge of state processing by efficiently combining the information from the graphs and the embedding.
@@ 10    $X_{H}^{(f)}(u):=v\textrm{ with }\forall u\in V_{H},v\in V_{P},u\in\phi^{(t+1)}(v)$
We now present a state transition algorithm that computes the next state based on the current state and action. This algorithm integrates the Node Embedding and Topology Adapting algorithms introduced in (Ngo et al., 2023a) as subroutines.
Algorithm 1 takes a current state $s_{t}$ at time step $t$ including a current logical graph $P$, a current hardware graph $H$, and current embedding $\phi^{(t)}$, as well as an action $a_{t}$ as input. It produces the next state $s_{t+1}$ including the logical graph $P$, the hardware graph $H$, and the next embedding $\phi^{(t+1)}$.
Before explaining this algorithm in detail, we recall the set of embedded nodes as $U^{(t)}_{P}$, which can be inferred from $\phi^{(t)}$. Algorithm 1 must guarantee that the resulting next embedding $\phi^{(t+1)}$ is a feasible embedding of $P[U^{(t)}_{P}\cup\{a_{t}\}]$ to $H$. To realize this, we first initialize the decision variable isolatedisolated as True (line 2) and assign $\phi$ as $\phi^{(t)}$ (line 3). Then, the node $a_{t}$ is embedded into the current hardware graph $H$ by finding an additional embedding $\phi^{\prime}$ through the Node Embedding algorithm. If the additional embedding $\phi^{\prime}$ is not empty ($a_{t}$ is successfully embedded), Algorithm 1 constructs a new embedding $\phi^{(t+1)}$ by combining $\phi$ and $\phi^{\prime}$ (line 7). In addition, it updates new features for the hardware graph $H$ following a hardware representation discussed in the previous section (line 8). The next state $s_{t+1}$ is a combination of logical graph $P$, hardware graph $H$ with updated features, and new embedding $\phi^{(t+1)}$. In the end, we assign the value of isolatedisolated to False, finish the loop, and return the new state $s_{t+1}$. Otherwise, if $\phi^{\prime}=\emptyset$, we cannot embed the node $a_{t}$ with the current embedding. We call this phenomenon the isolated problem. In order to overcome this bottleneck, we expand the current embedding $\phi$ using the Topology Adapting algorithm (line 11) and repeat the whole process until we find a satisfied additional embedding $\phi^{\prime}$.
From the lemmas provided in (Ngo et al., 2023a), it can be proven while assuming that the current embedding $\phi^{(t)}$ is feasible, the next embedding $\phi^{(t+1)}$ of the subsequent state $s_{t+1}$ returned by Algorithm 1 is a feasible embedding of $P[U^{(t)}_{P}\cup\{a\}]$ in $H$. Additionally, the empty embedding $\phi^{(0)}$ initialized at step 0 is a feasible embedding of the empty logical graph $P[\emptyset]$ in $H$. Consequently, considering $T=|V_{P}|$, we can recursively infer that the final embedding $\phi^{(T)}$ corresponding to the terminal state $s_{T}$ obtained in the $T$-th step is a feasible embedding of $P$ in $H$. Hence, with the state transition algorithm, CHARME effectively addresses the challenge of ensuring feasibility.
@@ 3.4. Exploration Strategy
Input: Set of generated logical graphs with $m$ elements $\mathcal{G}=\left(P_{1},\ldots,P_{m}\right)$, the hardware graph $H=(V_{H},E_{H})$, the set of baseline orders $\mathcal{\bar{O}}=\{\bar{O}_{1},\dots,\bar{O}_{m}\}$, the sampling limit $D$ and the exploration limit $K$
@@
2Initialize $\mathcal{O}=\left(O_{1},\ldots,O_{m}\right)$ where each $O_{j}$ is a random permutation of the set of nodes $V_{j}$ with $j=\{1,2,\dots,m\}$
3Initialize potential score $\mu_{j}$ as $|V_{H}|$, for each graph $P_{j}$ with $j=\{1,2,\dots,m\}$, and store in $\mathcal{M}=\left\{\mu_{1},\mu_{2},\mu_{3},\ldots,\mu_{m}\right\}$
@@
7 $O^{\prime},\_:=\texttt{Order Refining}(P_{i},H,0,\varnothing,\varnothing,K,F(O_{i}))$
8 if $O^{\prime}\neq\emptyset$ and $F(O^{\prime})<F(O_{i})$ then
@@ 9    $O_{i}\leftarrow O^{\prime}$
10    $\mu_{i}=F(O_{i})\div F(\bar{O}_{i})$
11
12
@@ 10    $\mu_{i}=F(O_{i})\div F(\bar{O}_{i})$
A key challenge our proposed RL model faces is maintaining efficiency when training on large instances. As the number of nodes in the training logical graphs increases, the number of possible action sequences grows exponentially. Consequently, it becomes more challenging for the model to explore good action sequences for learning, potentially leading to an inefficient policy. In CHARME, we address this issue by incorporating an exploration term in the reward function, as defined in Equation 9. The intuition behind this exploration term is to guide the model to replicate precomputed actions $\bar{a}_{t}$, during the initial training steps. As the policy ”matures” through updates, the model gradually reduces its reliance on precomputed actions and begins exploring by itself. This approach allows the model to start from a region with high-quality solutions and efficiently explores from there, enhancing the overall efficiency of training.
Precomputed actions can be extracted from the solutions of the embedding method in (Ngo et al., 2023a). However, that method focuses on providing a fast minor embedding, so the sequence of precomputed actions is not carefully refined. In this section, we propose a technique for exploring efficient sequences of precomputed actions in training logical graphs, which makes the training process more efficient. To avoid confusion here, we mention the sequences that include precomputed actions as embedding orders. We name our proposed technique as Order Exploration.
@@ 3.4.1. Overview
Given a logical graph $P=(V_{P},E_{P})$ with $T=|V_{P}|$ and a hardware graph $H=(V_{H},E_{H})$, an embedding order $O=(\bar{a}_{1},\dots,\bar{a}_{T})$ with $\bar{a}_{t}\in V_{P},t\in[1,T]$ results in a set of embeddings $\{\phi^{(1)},\dots,\phi^{(T)}\}$. The position of an action in $O$ is the step in which the action is embedded in the hardware graph, so we denote an index of actions in $O$ as an embedding step. As mentioned in previous sections, given an embedding step $t$, the embedding $\phi^{(t)}$ is constructed by inputting Algorithm 1 with $\phi^{(t-1)}$ and $\bar{a}_{t-1}$ for $t\in[1,T-1]$. We note that the construction of $\phi^{(t)}$ may require the topology expansion (i.e. the function Topology Adapting is triggered). We call embeddings that require topology expansion as expansion embeddings. Otherwise, we call them non-expansion embeddings. The expansion and non-expansion notions are used in the remainder of this section. In addition, we denote $F(O)=|\phi^{(T)}|$ as the efficiency score of $O$. Order Exploration aims to fast find embedding orders for training logical graphs such that their efficiency scores are as small as possible.
@@
The overview of Order Exploration is described in Algorithm 2. Algorithm 2 takes the inputs as the set of $m$ given training logical graphs, denoted as $\mathcal{G}=(P_{1},\ldots,P_{m})$, the set of baseline orders $\mathcal{\bar{O}}=\{\bar{O}_{1},\dots,\bar{O}_{m}\}$ (i.e., embedding orders of ATOM), the sampling limit $D$, and the exploration limit $K$. The training graph $P_{i}$ includes the set of nodes $V_{i}$ and the set of edges $E_{i}$. Given a logical graph $P_{i}\in\mathcal{G}$, we define the embedding order $O_{i}=(\bar{a}_{1},...,\bar{a}_{|V_{i}|})$ as the sequence of precomputed actions for the training graph $P_{i}$. The goal of Algorithm 2 is to quickly explore embedding orders for $m$ logical training graphs, denoted as $\mathcal{O}=(O_{1},\ldots,O_{m})$, such that their efficiency scores are as small as possible.
@@
In general, Algorithm 2 operates by randomly selecting a graph $P_{i}\in\mathcal{G}$ and exploring its order. Specifically, the algorithm begins by initializing the embedding orders for logical graphs in the training set $\mathcal{G}$ (line 1). Given a graph $P_{i}=(V_{i},E_{i})$, $O_{i}$ is initialized as a random permutation of the set $V_{i}$. Then, each graph $P_{i}$ is assigned an equal potential score $\mu_{i}$ (line 2), which reflects the potential for further improvement of its embedding order. Each potential score is initially set to the worst-case ratio between two possible solutions, which is given by $|V_{H}|$.
@@
Following initialization, the algorithm updates the embedding orders in $D$ steps. At each step, a graph $P_{i}$ is selected from the set $\mathcal{G}$ based on the distribution of potential scores (line 4). Specifically, the selection probability is determined as follows:
(13) $p\left(P_{i}\mid\mathcal{M}\right)=\frac{\mu_{i}}{\sum_{j}\mu_{j}}$
Afterward, the order $O_{i}$ is refined by the subroutine Order Refining (line 5) which is detailed in the next section as Algorithm 3. If the refined order $O^{\prime}$ is valid (i.e., $O^{\prime}\neq\emptyset$) and achieves a better score than the current order $O_{i}$ (i.e., $F(O^{\prime})<F(O_{i})$), we re-assign $O_{i}$ as $O^{\prime}$ (line 7), and update the potential score $\mu_{i}$ as follows:
(14) $\mu_{i}=\frac{F(O_{i})}{F(\bar{O}_{i})}$
@@
We recall that $F(\bar{O}_{i})$ is the total number of qubits used for embedding $P_{i}$ resulting from ATOM (Ngo et al., 2023a). Thus, the score $\mu_{i}$, calculated by Equation 14, gives us the gap between our refined solution and the existing solution. Consequently, the larger the gap, the higher the likelihood that the corresponding graph will be selected for further refinement. It is also worth noting that if the initial potential scores are sufficiently large, every graph can be selected at least once. When the algorithm finishes discovering, it returns all feasible orders through the set $\mathcal{O}$ (line 9).